# Adaptive Reduced Multilevel Splitting

### v1

@@ 4.1 AMS
We wish to draw a sample of size $N$ according to the importance sampling distribution $\mu^{(k)}_{l^{(k)}}$ in the form (1.3), where, for ease of presentation, we temporarily assume that the target level ${l^{(k)}}$ is known. We suppose we start with a given sample of size $N$ aproximately distributed according to the law $\mu^{(k)}_{l_{0}}$ with $l_{0}<{l^{(k)}}$, as well as with a given estimator of the associated normalization $Z^{(k)}_{l_{0}}$. The AMS algorithm [5] will be used here to generate a finite sequence of $N$-samples approximately distributed according to a sequence of nested distributions $\{\mu^{(k)}_{l_{i}}\}_{i\geq 1}$, as well as estimators of the associated normalizations $\left\{Z^{(k)}_{l_{i}}\right\}_{i\geq 1}$, i.e., such that $l_{0}<l_{1}<\ldots<l_{m}=l^{(k)}$. At iteration $i\geq 0$, the empirical distribution related to the $N$-sample $(\Xi^{(i)}_{1},\ldots,\Xi^{(i)}_{N})$ (we omit in notation the dependence in $(k)$ ) generated by AMS will be denoted
@@
Let us describe briefly the AMS algorithm. At its $i$-th iteration, the algorithm performs sucessively: i) a weighting and selection step, and ii) a mutation step. The weighting step i) assigns weights to the empirical distribution ${\mu^{N,{(k)}}_{l_{i}}}$ so that it approximates $\mu_{l_{i+1}}^{(k)}$, up to a normalization constant given by an empirical approximation of the normalization ratios. The details work as follows. The empirical distribution at level $l_{i+1}$ and the associated normalizations are obtained by weighting of the current $N$-sample:
@@
where the last equality is well-defined because $l_{i}<l_{i+1}$. This empirical distribution of course requires the determination (or a choice) of $l_{i+1}$. AMS performs an adaptive choice of $l_{i+1}$ as follows: AMS starts by sorting the $N$ particles according to their scores in ascending order. AMS then considers an adaptive target distribution defined by a fixed number $M<N$ of particles. The latter are assigned a weight equal to zeroa, and the level ${l_{i+1}}$ is defined by the $M$-th smallest particle score [^4] To ease the presentation we assume here that particles have different scores. A good practice in case of equalities is to discard all the particles with the same score, and locally change the precise value of $M$ accordingly.
@@
with the quantile (proportion of killed particles) satisfies $M=\theta N$.
In the selection step, the first $M$ particles with zero-weights are removed and surviving particles are duplicated. The $M$ duplicated particles are chosen uniformly among the surviving particles. The AMS approximation of the normalization constant thus follows the recursion
@@
${Z^{N,{(k)}}_{l^{(k)}}}={Z^{N,{(k)}}_{l_{m}}}={Z^{N,{(k)}}_{l_{0}}}(1-{M}/{N})^{m}.$
The mutation step ii) then consists of a Markovian transition that leaves $\mu_{l_{i+1}}^{(k)}$ invariant and is applied individually and independently $M$ times to the $M$ duplicated particles, the goal being to bring diversity in the $N$-sample to counterbalance the effect of selection. In practice, Markovian transitions are performed using a Metropolis-Hasting algorithm. We mention that the Markovian kernel is often chosen to be adaptive in order to maintain a given Metropolis acceptance rate [14].
@@ 4.2 Reaching the critical level with AMS
Let us now describe the main Monte Carlo sampling routine of the algorithm. We recall that after $k$ iterations, one is given with a current reduced score $S^{(k)}$, an associated pointwise error estimate $E^{(k)}$, an initial level $l_{\mathrm{bridge}}(k)$ and a sample of size $N$ approximately distributed according to $\mu_{l_{\mathrm{bridge}}(k)}^{(k)}$, that we dnote using a sample-size superscipt: $\mu_{l_{\mathrm{bridge}}(k)}^{N,(k)}$ One want to perform an AMS simulation targeting the distribution $\mu_{l^{(k)}}^{(k)}$, where we recall that ${l^{(k)}}$ is the critical, non-achievable level the closest to $l_{\mathrm{bridge}}(k)$. The latter is formally defined by (2.2) (for $N=+\infty$). When using sample-based empirical distributions, one will need to adapt the latter conditions.
The AMS simulation will proceed as explained in Section 4.1, creating a sequence $(l_{i})_{i\geq 0}$ of increasing levels starting from $l_{0}={l_{\mathrm{bridge}}(k)}$, creating a sequence of empirical distributions ${\mu^{N,{(k)}}_{l_{i}}}$ for $i\geq 0$. We then check for each proposed level increase from $l_{i}$ to $l_{i+1}$ that the entropic conditions (2.2) using the empirical distribution ${\mu^{N,{(k)}}_{l_{i}}}$ to approximate the latter. More precisely, we remark that
@@
$\min\{l_{i},\,i\geq 0:\;\ln\frac{{\mu^{N,{(k)}}_{l_{i}}}(\,{\rm l}{\rm l}_{S^{(k)}>l_{i+1}})}{{\mu^{N,{(k)}}_{l_{i}}}(\,{\rm l}{\rm l}_{S^{(k)}-E^{(k)}>l_{i+1}})}\geq e\;\;\text{or}\;\;\ln\frac{{\mu^{N,{(k)}}_{l_{i}}}(\,{\rm l}{\rm l}_{S^{(k)}>l_{i+1}})}{{\mu^{N,{(k)}}_{l_{i}}}(\,{\rm l}{\rm l}_{S^{(k)}+E^{(k)}>{l_{\max}}})}=+\infty\}.$
@@
If the logarithm of the ratio in (4.3) is finite when $N\to\infty$ for all intermediate levels from $\infty$- ∞ to $l_{i}$, for each of the AMS intermediate levels $l_{j}$ with $j\leq i$ then we can verify recursively that, at a formal level,
@@
Algorithm 2 (Simulation up to the critical level)
1:level of interest $l_{\mathrm{max}}$, initial distribution $\mu^{N,(k)}_{l_{0}(k)}$ and normalization $Z^{N,(k)}_{l_{0}(k)}$, related reduced score $S^{(k)}$ and error $E^{(k)}$, quantile $\theta\in[1/N,1[$, worst-case log cost threshold $e$.
@@
3:Set $i=0$, $e_{try}=0$, $M=\theta N\geq 1$
4:while  $l_{i}<l_{\mathrm{max}}$ $e_{try}\leq e$ $d_{try}<+\infty$  do
@@
$e_{try}=\ln\frac{{\mu^{N,{(k)}}_{l_{i}}}(\,{\rm l}{\rm l}_{S^{(k)}>l_{i+1}})}{{\mu^{N,{(k)}}_{l_{i}}}(\,{\rm l}{\rm l}_{S^{(k)}-E^{(k)}>l_{i+1}})}\simeq\mathrm{Ent}\left(\check{\eta}_{l_{i+1}}^{(k)}\mid\mu_{l_{i+1}}^{(k)}\right)$
@@
8:    if  $l_{i+1}<l_{\mathrm{max}}$ $e_{try}\leq e$ $d_{try}<+\infty$ then
9:         Update ${Z^{N,{(k)}}_{l_{i+1}}}$ with recursion (4.1)
10:         Kill the $M$ particles with the lowest score
@@
The following remark is useful in practice, as it provides some prior information on the worst-case log cost threshold parameter $e$.
@@
Consider, in Algorithm $2$ above, the empirical distribution ${\mu_{l_{i}}^{N,(k)}}$. By construction of AMS, $l_{i+1}$ is the $M$-th smallest score among the $N$ scores. Set $e_{\min}(M,N)\displaystyle\mathop{=}^{{\rm def}}\ln\frac{N-M}{N-M-1}$. As long as $E^{(k)}$ is non zero, at least one particle of the sample, the $M$-th particle, lies outside the level set obtained by subtracting the error to the score. $e_{\min}(M,N)$ is thus the lower bound of the entropy between the empirical measures with and without the error term, that is
$e_{\min}(M,N)\leq\ln\frac{\mu_{l_{i}}^{N,(k)}(\,{\rm l}{\rm l}_{S^{(k)}>l_{i+1}})}{\mu_{l_{i}}^{N,(k)}(\,{\rm l}{\rm l}_{S^{(k)}-E^{(k)}>l_{i+1}})}.$
Setting the numerical threshold to the minimal possible value $e=e_{\min}(M,N)$ can be considered a safe strategy, leading to a slower but safer convergence of the algorithm.
@@ 4.3 Bridging with an AMS step
As exposed in Section 2.4, the bridging procedure at the end of iteration $k$ consist in selecting a past index $=k_{\mathrm{bridge}}:=k_{\mathrm{bridge}}(k)\leq k$ and a new initial level ${l_{\mathrm{bridge}}(k+1)}$ which will be used in the AMS simulation with the new score approximation $S^{(k+1)}$ starting from the past, recorded, empirical distribution $\mu^{N,{(k_{\mathrm{bridge}})}}_{l^{(k_{\mathrm{bridge}})}}$. As defined previously, bridging consists in choosing, using a lexicographic search, a feasible pair $(k_{\mathrm{bridge}},{l_{\mathrm{bridge}}(k+1)})\in\llbracket 0,k\rrbracket\times\mathbb{R}$, where feasibility means satisfying the quantile condition (2.5) and the two entropic conditions  (2.6) and (2.7). It is now necessary to define empirical counterparts to the latter conditions. As previously, the two entropic conditions are approximated by
$\mathrm{Ent}\left(\check{\eta}_{l^{\prime}}^{(k+1)}\mid\mu_{l^{\prime}}^{(k+1)}\right)\simeq\ln\frac{{\mu^{N,{(k^{\prime})}}_{l^{(k^{\prime})}}}(\,{\rm l}{\rm l}_{S^{(k+1)}>{l^{\prime}}})}{{\mu^{N,{(k^{\prime})}}_{l^{(k^{\prime})}}}(\,{\rm l}{\rm l}_{S^{(k+1)}-E^{(k+1)}>{l^{\prime}}})}\leq e,$ (4.4)
@@
where in the above $(k^{\prime},l^{\prime})$ denote feasible pairs that are candidates for the solution $(k_{\mathrm{bridge}},l_{\mathrm{bridge}}(k+1))$. Note that the well-posedness of (4.5) can be justified by Remark 4.1. Concerning the quantile condition, we remark that, given some $k^{\prime}$ and ${l^{\prime}}$, and assuming that the domination relation stands
@@
In order to verify (4.6), we propose to browse the history of past AMS simulations and check the inclusion of the support of $\mu_{l^{\prime}}^{(k+1)}$ in the one of ${\mu_{l^{(k^{\prime})}}^{(k^{\prime})}}$ using all the recorded N-samples targeting the past proposal distributions ${\mu_{l^{(\tilde{k})}}^{N,(\tilde{k})}}$ with $\tilde{k}<k_{\mathrm{bridge}}$, whose samples are scattered between the reference distribution and the current distribution. This check must be carried out for all $\tilde{k}<k_{\mathrm{bridge}}$ and is given by
@@
To sum up, bridging consist in selecting some feasible pair, i.e., some pair $(k_{\mathrm{bridge}},{l_{\mathrm{bridge}}(k+1)})$ satisfying (4.4), (4.5), (4.7) and (4.8). Note that once the pair has been determined, the normalization estimate can be updated as
${Z^{N,{(k+1)}}_{l_{\mathrm{bridge}}(k+1)}}={Z^{N,{({k_{\mathrm{bridge}}})}}_{l^{(k_{\mathrm{bridge}})}}}{\mu_{l^{(k_{\mathrm{bridge}})}}^{N,{({k_{\mathrm{bridge}}})}}}(\,{\rm l}{\rm l}_{S^{(k+1)}\geq{l_{\mathrm{bridge}}(k+1)}}).$ (4.9)
Then the $\tilde{M}(k_{\mathrm{bridge}})$ particles (at max given by $\tilde{M}(k_{\mathrm{bridge}})=\theta N$ when (4.7) is saturated) with updated scores below or equal to ${l_{\mathrm{bridge}}(k+1)}$ are replaced by duplicating surviving particles and performing mutation targeting $\mu_{l_{\mathrm{bridge}}(k+1)}^{(k+1)}$ (in an AMS line of thinking). The next AMS simulation with $S^{(k+1)}$ can then be started from the level ${l_{\mathrm{bridge}}(k+1)}$. The following iteration then starts up to the critical level $l^{(k+1)}=l^{(k+1)}(e,{l_{\mathrm{bridge}}(k+1)})$, an so on.
As mentioned in Section 2.4, among the feasible set, we focus on the following two pairs.
  • There exists the trivial pair
    $(k_{\mathrm{bridge}},{l_{\mathrm{bridge}}(k+1)})=(0,-\infty).$
    This trivial choice restarts the AMS simulation with the new score $S^{(k+1)}$ from the beginning. However, as mentioned before, this option entails a significant computational overhead if the complexity of evaluating the reduced score is not negligible.
  • An alternative is to search for the largest (lexicographic) pair $(k_{\mathrm{bridge}},{l_{\mathrm{bridge}}(k+1)})\in\llbracket 0,k\rrbracket\times\mathbb{R}$. A simple way to solve this maximisation problem is to begin with the highest possible index $k^{\prime}:=k$ and then to decrease $k^{\prime}$ by a unit if no feasible level $l^{\prime}$ has been found for this index. To search for the largest feasible level $l^{\prime}$ at a given index $k^{\prime}$, one can proceed as follows. Begin by setting $l^{\prime}$ to the highest level satisfying (4.7), then lower this level until (4.4) and (4.5) hold and finally perform for all $\tilde{k}<k_{\mathrm{bridge}}$ the check (4.8).
The bridging procedure is summarized by the following algorithm.
@@
1:level of interest $l_{\mathrm{max}}$, current index $k$, for all $k^{\prime}\leq k$, particles of distribution $\mu^{N,(k^{\prime})}_{l^{(k^{\prime})}}$, normalization $Z^{N,(k^{\prime})}_{l^{(k^{\prime})}}$ and related reduced score $S^{(k^{\prime})}$ and error $E^{(k^{\prime})}$, reduced score $S^{(k+1)}$, the quantile $\theta=M/N\in(0,1)$, worst-case log cost threshold $e$.
@@
6:    Set $\tilde{M}=N\theta$
@@
9:         Compute (4.4): $e_{try}=\ln\frac{{\mu^{N,{(k^{\prime})}}_{l^{(k^{\prime})}}}(\,{\rm l}{\rm l}_{S^{(k+1)}>{l^{\prime}}})}{{\mu^{N,{(k^{\prime})}}_{l^{(k^{\prime})}}}(\,{\rm l}{\rm l}_{S^{(k+1)}-E^{(k+1)}>{l^{\prime}}})}\simeq\mathrm{Ent}\left(\check{\eta}_{l^{\prime}}^{(k+1)}\mid\mu_{l^{\prime}}^{(k+1)}\right)$
10:         Compute (4.5): * $d_{try}=\ln\frac{{\mu^{N,{(k^{\prime})}}_{l^{(k^{\prime})}}}(\,{\rm l}{\rm l}_{S^{(k+1)}>{l^{\prime}}})}{{\mu^{N,{(k^{\prime})}}_{l^{(k^{\prime})}}}(\,{\rm l}{\rm l}_{S^{(k+1)}+E^{(k+1)}>{l_{\max}}})}\simeq\mathrm{Ent}\left(\hat{\eta}_{l_{\max}}^{(k+1)}\mid\mu_{l^{\prime}}^{(k+1)}\right)$
11:         if $e_{try}\leq e\,\&\,d_{try}<+\infty$ then
@@
24:    Set the largest pair as $(k_{\mathrm{bridge}},l_{\mathrm{bridge}}(k+1))=(k^{\prime},l^{\prime})$
25:    Compute ${Z^{N,{(k+1)}}_{l_{\mathrm{bridge}}(k+1)}}$ using (4.9)
26:    Kill the $\tilde{M}$ particles in $\mu_{l^{(k_{\mathrm{bridge}})}}^{N,(k_{\mathrm{bridge}})}$ of smallest score $S^{(k+1)}$
27:    Perform $\tilde{M}$ duplications (uniformly among the survivors)
28:    Simulate $\mu_{l_{\mathrm{bridge}}(k+1)}^{N,(k+1)}$ by mutating the duplicates to target $\mu_{l_{\mathrm{bridge}}(k+1)}^{(k+1)}$
@@
30:    Set $l_{\mathrm{bridge}}(k+1)=-\infty$
31:    Set ${Z^{N,{(k+1)}}_{l_{\mathrm{bridge}}(k+1)}}=1$
32:    Simulate $\mu_{l_{\mathrm{bridge}}(k+1)}^{N,(k+1)}$ (by sampling the reference distribution $\pi$)
@@
34:Return Empirical measure $\mu_{l_{\mathrm{bridge}}(k+1)}^{N,(k+1)}$ and normalizing constant ${Z^{N,{(k+1)}}_{l_{\mathrm{bridge}}(k+1)}}$

### v3

@@ 3 Practical Implementation: AMS Simulations
In this section, we propose a practical implementation of the idealized algorithm introduced in Section 2.7 using an AMS methodology. The latter will simulate the flow of proposals for growing level parameters $l$ by providing Monte Carlo samples with approximately correct distributions for large sample sizes.
@@ 3.1 AMS
We wish to draw a sample of size $N$ according to the importance sampling distribution $\mu^{(k)}_{l^{(k)}}$ in the form (1.3), where, for ease of presentation, we temporarily assume that the target level ${l^{(k)}}$ is known. We suppose we start with a given sample of size $N$ aproximately distributed according to the law $\mu^{(k)}_{l_{b}}$ with $l_{b}<{l^{(k)}}$, as well as with a given estimator of the associated normalization $Z^{(k)}_{l_{b}}$. The AMS algorithm [5] will be used here to generate a finite sequence of $N$-samples approximately distributed according to a sequence of nested distributions $\{\mu^{(k)}_{l_{i}}\}_{i\geq 1}$, as well as estimators of the associated normalizations $\left\{Z^{(k)}_{l_{i}}\right\}_{i\geq 1}$, i.e., such that $l_{b}<l_{1}<\ldots<l_{m}=l^{(k)}$. At iteration $i\geq 0$, the $N$-sample of particles (clones) will be denoted
$(\Xi^{(i)}_{1},\ldots,\Xi^{(i)}_{N})$
and we omit in notation the dependence in $(k)$. The associated empirical distribution will be denoted
@@
Let us describe briefly the AMS algorithm. At the $i$-th iteration, the algorithm performs sucessively: i) a weighting and selection step, and ii) a mutation step. The weighting step i) assigns weights to the empirical distribution ${\mu^{N,{(k)}}_{l_{i}}}$ so that it approximates $\mu_{l_{i+1}}^{(k)}$, up to a normalization constant given by an empirical approximation of the normalization ratios. More precisely, the empirical distribution at level $l_{i+1}$ and the associated normalizations are obtained by weighting the current $N$-sample:
@@
where the equality is well-defined because $l_{i}<l_{i+1}$. This empirical distribution of course requires the determination (or a choice) of $l_{i+1}$. AMS performs an adaptive choice of $l_{i+1}$ as follows. AMS starts by sorting the $N$ particles according to their scores in ascending order. AMS then considers an adaptive target distribution defined by a fixed number $M<N$ of particles. The latter are assigned a weight equal to zero, and the level ${l_{i+1}}$ is defined by the $M$-th smallest particle score
@@
with the quantile (proportion of killed particles) satisfies $M=\lfloor\theta N\rfloor$.
In the selection step, the first $M$ particles with zero-weights are removed and surviving particles are duplicated. The $M$ resampled particles are chosen uniformly among the surviving particles. The AMS approximation of the normalization constant thus follows the recursion [^3] To ease the presentation we have assumed here that particles have different scores. However, in the case $\exists\mathcal{J}\subset\mathbb{N}$ such that $S^{(k)}(\Xi^{(i)}_{(M)})=S^{(k)}(\Xi^{(i)}_{(M+j)})$ for $j\in\mathcal{J}$, a good practice is to kill in addition the set of particles $\{S^{(k)}(\Xi^{(i)}_{(M+j)})\}_{j\in\mathcal{J}}$, and locally change the value of $M$ in (3.1) to $M+\textrm{card}(\mathcal{J})$ accordingly.
@@
${Z^{N,{(k)}}_{l^{(k)}}}={Z^{N,{(k)}}_{l_{m}}}={Z^{N,{(k)}}_{l_{b}}}(1-{M}/{N})^{m}.$
The mutation step ii) then consists of a Markovian transition that leaves $\mu_{l_{i+1}}^{(k)}$ invariant and is applied individually and independently a prescribed number of times to the $M$ particles after resampling, the goal being to bring diversity in the $N$-sample to counterbalance the effect of selection. In practice, Markovian transitions can be performed using a Metropolis-Hasting algorithm or other techniques studied in the Markov Chain Monte Carlo (MCMC) literature. We mention that the Markovian kernel is often chosen to be adaptive in order to maintain a given Metropolis acceptance rate [15].
@@ 3.2 Reaching the critical level with AMS
Let us now describe the main Monte Carlo sampling routine of the algorithm. We recall that after $k$ iterations, one is given with a current reduced score $S^{(k)}$, an associated pointwise error estimate $E^{(k)}$, an initial level $l_{b}(k)$ and a sample of size $N$ approximately distributed according to $\mu_{l_{b}(k)}^{(k)}$, that we denote using a sample-size superscript: $\mu_{l_{b}(k)}^{N,(k)}$ One want to perform an AMS simulation targeting the distribution $\mu_{l^{(k)}}^{(k)}$, where we recall that ${l^{(k)}}$ is the critical, non-achievable level the closest to $l_{b}(k)$. The latter is formally defined by (2.2) (for $N=+\infty$). When using sample-based empirical distributions, one will need to adapt the latter conditions.
The AMS simulation will proceed as explained in Section 3.1, creating a sequence $(l_{i})_{i\geq 0}$ of increasing levels starting from $l_{b}={l_{b}(k)}$, creating a sequence of empirical distributions ${\mu^{N,{(k)}}_{l_{i}}}$ for $i\geq 0$. We then check for each proposed level increase from $l_{i}$ to $l_{i+1}$ that the entropic conditions (2.2) using the empirical distribution ${\mu^{N,{(k)}}_{l_{i}}}$ to approximate the latter. More precisely, we remark that
@@
$\min\{l_{i},\,i\geq 0:\;\ln\frac{{\mu^{N,{(k)}}_{l_{i}}}(\,{\rm l}{\rm l}_{S^{(k)}>l_{i+1}})}{{\mu^{N,{(k)}}_{l_{i}}}(\,{\rm l}{\rm l}_{S^{(k)}-E^{(k)}>l_{i+1}})}\geq c\;\;\text{or}\;\;\ln\frac{{\mu^{N,{(k)}}_{l_{i}}}(\,{\rm l}{\rm l}_{S^{(k)}>l_{i+1}})}{{\mu^{N,{(k)}}_{l_{i}}}(\,{\rm l}{\rm l}_{S^{(k)}+E^{(k)}>{l_{\max}}})}=+\infty\}.$
@@
If the logarithm of the ratio in (3.3) is finite when $N\to\infty$ for all intermediate levels $l_{j}$ from $\infty$- ∞ to $l_{i}$then we can verify recursively that, at a formal level,
@@
Algorithm 2 (AMS simulation up to the critical level)
1:level of interest $l_{\mathrm{max}}$, initial distribution $\mu^{N,(k)}_{l_{b}(k)}$ and normalization $Z^{N,(k)}_{l_{b}(k)}$, related reduced score $S^{(k)}$ and error $E^{(k)}$, quantile $\theta\in[1/N,1[$, worst-case log cost threshold $c$.
@@
3:Set $i=0$, $c_{try}=0$
4:while  $l_{i}<l_{\mathrm{max}}$ $c_{try}\leq c$ $d_{try}<+\infty$  do
@@
$c_{try}=\ln\frac{{\mu^{N,{(k)}}_{l_{i}}}(\,{\rm l}{\rm l}_{S^{(k)}>l_{i+1}})}{{\mu^{N,{(k)}}_{l_{i}}}(\,{\rm l}{\rm l}_{S^{(k)}-E^{(k)}>l_{i+1}})}\simeq\mathrm{Ent}\left(\check{\eta}_{l_{i+1}}^{(k)}\mid\mu_{l_{i+1}}^{(k)}\right)$
@@
8:    if  $l_{i+1}<l_{\mathrm{max}}$ $c_{try}\leq c$ $d_{try}<+\infty$ then
9:         Kill the $M\geq\lfloor\theta N\rfloor$ particles with the lowest score or with score equal to $l_{i+1}$
10:         Update ${Z^{N,{(k)}}_{l_{i+1}}}$ with recursion (3.1)
@@
The following remark is useful in practice, as it provides some prior information on the worst-case log cost threshold parameter $c$.
@@
Consider, in Algorithm $2$ above, the empirical distribution ${\mu_{l_{i}}^{N,(k)}}$. By construction of AMS, $l_{i+1}$ is the $M$-th smallest score among the $N$ scores. Set $c_{\min}(M,N)\displaystyle\mathop{=}^{{\rm def}}\ln\frac{N-M}{N-M-1}$. As long as $E^{(k)}$ is non zero, at least one particle of the sample, the $M$-th particle, lies outside the level set obtained by subtracting the error to the score. $c_{\min}(M,N)$ is thus the lower bound of the entropy between the empirical measures with and without the error term, that is
$c_{\min}(M,N)\leq\ln\frac{\mu_{l_{i}}^{N,(k)}(\,{\rm l}{\rm l}_{S^{(k)}>l_{i+1}})}{\mu_{l_{i}}^{N,(k)}(\,{\rm l}{\rm l}_{S^{(k)}-E^{(k)}>l_{i+1}})}.$
Setting the numerical threshold to the minimal possible value $c=c_{\min}(M,N)$ can be considered a safe strategy, leading to a slower but safer convergence of the algorithm.
@@ 3.3 Bridging with an AMS step
As exposed in Section 2.4, the bridging procedure at the end of iteration $k$ consist in selecting a past index $=k_{b}:=k_{b}(k)\leq k$ and a new initial level ${l_{b}(k+1)}$ which will be used in the AMS simulation with the new score approximation $S^{(k+1)}$ starting from a past (appropriately recorded) empirical distribution $\mu^{N,{(k_{b})}}_{l^{(k_{b})}}$. As defined previously, bridging consists in choosing, using a lexicographic search, a feasible pair $(k_{b},{l_{b}(k+1)})\in\llbracket 0,k\rrbracket\times\mathbb{R}$, where feasibility means satisfying the quantile condition (2.5) and the two entropic conditions  (2.6) and (2.7). It is now necessary to define empirical counterparts to the latter conditions. As previously, the two entropic conditions are approximated by
$\mathrm{Ent}\left(\check{\eta}_{l^{\prime}}^{(k+1)}\mid\mu_{l^{\prime}}^{(k+1)}\right)\simeq\ln\frac{{\mu^{N,{(k^{\prime})}}_{l^{(k^{\prime})}}}(\,{\rm l}{\rm l}_{S^{(k+1)}>{l^{\prime}}})}{{\mu^{N,{(k^{\prime})}}_{l^{(k^{\prime})}}}(\,{\rm l}{\rm l}_{S^{(k+1)}-E^{(k+1)}>{l^{\prime}}})}\leq c,$ (3.4)
@@
where in the above $(k^{\prime},l^{\prime})$ denote feasible pairs that are candidates for the solution $(k_{b},l_{b}(k+1))$. Note that the well-posedness of (3.5) can be justified by Remark 3.1. Concerning the quantile condition, we remark that, given some $k^{\prime}$ and ${l^{\prime}}$, and assuming that the domination relation stands
@@
In order to verify (3.6), we propose to browse the history of past AMS simulations and check the inclusion of the support of $\mu_{l^{\prime}}^{(k+1)}$ in the one of ${\mu_{l^{(k^{\prime})}}^{(k^{\prime})}}$ using all the recorded $N$-samples targeting the past proposal distributions ${\mu_{l^{(\tilde{k})}}^{N,(\tilde{k})}}$ with $\tilde{k}<k_{b}$, whose samples are scattered between the reference distribution and the current distribution. This check must be carried out for all $\tilde{k}<k_{b}$ and is given by
@@
To sum up, bridging consist in selecting some feasible pair, i.e., some pair $(k_{b},{l_{b}(k+1)})$ satisfying (3.4), (3.5), (3.7) and (3.8). Note that once the pair has been determined, the normalization estimate can be updated as
${Z^{N,{(k+1)}}_{l_{b}(k+1)}}={Z^{N,{({k_{b}})}}_{l^{(k_{b})}}}{\mu_{l^{(k_{b})}}^{N,{({k_{b}})}}}(\,{\rm l}{\rm l}_{S^{(k+1)}>{l_{b}(k+1)}}).$ (3.9)
Then the $\tilde{M}(k_{b})$ particles (at max given by $\tilde{M}(k_{b})=\lfloor\theta N\rfloor$ when (3.7) is saturated) with updated scores below or equal to ${l_{b}(k+1)}$ are replaced by duplicating surviving particles and performing mutation targeting $\mu_{l_{b}(k+1)}^{(k+1)}$ (in an AMS line of thinking). The next AMS simulation with $S^{(k+1)}$ can then be started from the level ${l_{b}(k+1)}$. The following iteration then starts up to the critical level $l^{(k+1)}(c,{l_{b}(k+1)})$, an so on.
As mentioned in Section 2.4, among the feasible set, we focus on either the trivial pair $(k_{b},{l_{b}(k+1)})=(0,-\infty)$, or the largest (lexicographic) pair $(k_{b},{l_{b}(k+1)})\in\llbracket 0,k\rrbracket\times\mathbb{R}$. A simple way to solve this maximisation problem is to begin with the highest possible index $k^{\prime}:=k$ and then to decrease $k^{\prime}$ by a unit if no feasible level $l^{\prime}$ has been found for this index. To search for the largest feasible level $l^{\prime}$ at a given index $k^{\prime}$, one can proceed as follows. Begin by setting $l^{\prime}$ to the highest level satisfying (3.7), then lower this level until (3.4) and (3.5) hold and finally perform for all $\tilde{k}<k_{b}$ the check (3.8).
The non-trivial bridging procedure is summarized by the following algorithm.
@@
1:level of interest $l_{\mathrm{max}}$, current index $k$, for all $k^{\prime}\leq k$, particles of distribution $\mu^{N,(k^{\prime})}_{l^{(k^{\prime})}}$, normalization $Z^{N,(k^{\prime})}_{l^{(k^{\prime})}}$ and related reduced score $S^{(k^{\prime})}$ and error $E^{(k^{\prime})}$, reduced score $S^{(k+1)}$, the quantile $\theta\in[1/N,1[$, worst-case log cost threshold $c$.
@@
6:    Set $\tilde{M}=\lfloor N\theta\rfloor$
@@
9:         Compute (3.4): $c_{try}=\ln\frac{{\mu^{N,{(k^{\prime})}}_{l^{(k^{\prime})}}}(\,{\rm l}{\rm l}_{S^{(k+1)}>{l^{\prime}}})}{{\mu^{N,{(k^{\prime})}}_{l^{(k^{\prime})}}}(\,{\rm l}{\rm l}_{S^{(k+1)}-E^{(k+1)}>{l^{\prime}}})}\simeq\mathrm{Ent}\left(\check{\eta}_{l^{\prime}}^{(k+1)}\mid\mu_{l^{\prime}}^{(k+1)}\right)$
10:         Compute (3.5): $d_{try}=\ln\frac{{\mu^{N,{(k^{\prime})}}_{l^{(k^{\prime})}}}(\,{\rm l}{\rm l}_{S^{(k+1)}>{l^{\prime}}})}{{\mu^{N,{(k^{\prime})}}_{l^{(k^{\prime})}}}(\,{\rm l}{\rm l}_{S^{(k+1)}+E^{(k+1)}>{l_{\max}}})}\simeq\mathrm{Ent}\left(\hat{\eta}_{l_{\max}}^{(k+1)}\mid\mu_{l^{\prime}}^{(k+1)}\right)$
11:         if $c_{try}\leq c\,\&\,d_{try}<+\infty$ then
@@
24:    Set the largest pair as $(k_{b},l_{b}(k+1))=(k^{\prime},l^{\prime})$
25:    Compute ${Z^{N,{(k+1)}}_{l_{b}(k+1)}}$ using (3.9)
26:    Kill the $\tilde{M}$ particles in $\mu_{l^{(k_{b})}}^{N,(k_{b})}$ of smallest score $S^{(k+1)}$ or with score equal to $l_{b}(k+1)$
27:    Perform duplications (uniformly among the survivors) to replace the killed particles
28:    Simulate $\mu_{l_{b}(k+1)}^{N,(k+1)}$ by mutating the duplicates to target $\mu_{l_{b}(k+1)}^{(k+1)}$
@@
30:    Set $l_{b}(k+1)=-\infty$
31:    Set ${Z^{N,{(k+1)}}_{l_{b}(k+1)}}=1$
32:    Simulate $\mu_{l_{b}(k+1)}^{N,(k+1)}$ (by sampling the reference distribution $\pi$)
@@
34:Return Empirical measure $\mu_{l_{b}(k+1)}^{N,(k+1)}$ and normalizing constant ${Z^{N,{(k+1)}}_{l_{b}(k+1)}}$